If a map drawn to the scale of 1.25 centimeters equals 50 kilometers, what distance does 3.5 centimeters represent?
Answer by checkley79(3341)

1.25/50=3.5/X cross multiply.
1.25x=3.5*50
1.25x=175
x=175/1.25
x=140 cm. ans.